OpenStudy (anonymous):

Integrate [0, 1/2] xcos(pi*x) Every method by parts I've tried so far has turned out wrong.

OpenStudy (richyw):

Integration by parts should work fine. set \(g=x\) and \(dh=\cos(\pi x)dx\)

OpenStudy (richyw):

This for sure works, I just did it. Using integration by parts and choosing what I had above. Then just use a simple U-substitution.

OpenStudy (anonymous):

[Integrate [0, 1/2] xcos(pi*x)=(pi-2)/2(pi)^2

OpenStudy (richyw):

well yeah you can plug it into a calculator, haha. you can also plug it into wolfram. this integration is easy enough I'm almost positive it will show you the correct steps.

OpenStudy (anonymous):

[Integrate [0, 1/2] xcos(pi*x let u=x so that du=dx and v=intgral cos (xpi)dx v=(1/pi)sin(pi*x) integration by parts uv-itgral[0,1/2]vdu just plug ins (1/pi)sinpi*x]-(1/pi)itgrlsin(pi*x)dx from 0 to 1/2 (1/pi)x sinpi*x - (1/pi)[-(1/pi) cos pi*x] from 0 to 1/2 =(1/2pi)+(1/pi^2)[cos pi*x/2-cos 0] =1/2pi - 1/2pi^2 =(pi-2)/2pi^2 ans